OpenStudy (anonymous):

Which of the following is not a question to ask when determining whether an expression involving exponents is completely simplified? A. Are all parentheses gone? B. Are all exponents positive? C. Have all negative constants been eliminated? D. Have all powers with the same base been combined?

OpenStudy (ribhu):

D

OpenStudy (mathmate):

Hints: Examples of what you would see if you follow the given steps A. Are all parentheses gone? \((2*3)^4=6^4\) B. Are all exponents positive? \(4^{-3}=\frac{1}{4^3}\) C. Have all negative constants been eliminated? \((-5)\times 3^4=5log\frac{1}{e}\times 3^4\) D. Have all powers with the same base been combined? \(4^6\times 2^{-4}=2^{12-4}=2^8\) Decide on which are reasonable steps to simplify an exponential expression (and which is not!)

OpenStudy (anonymous):

im done already